OpenStudy (howard-wolowitz):

Just two hard questions:

OpenStudy (xapproachesinfinity):

pufff dislike lengthy problems

OpenStudy (anonymous):

would you like help with the first one

OpenStudy (howard-wolowitz):

yes please

OpenStudy (howard-wolowitz):

alright i gotcha so far

OpenStudy (anonymous):

chair A requires 1 hour of waxing chair B requires 4 hours of waxing The maximum number of hours available (per week) for waxing is 100 hours. If x = number of chair A produced, y = number of chair B produced, we have: 1x + 4y <= 100 Similarly chair A requires 5 hour of painting chair B requires 6 hours of painting The maximum number of hours available (per week) for painting is 60 hours. If x = number of chair A produced, y = number of chair B produced, we have: 5x + 6y <= 60

OpenStudy (howard-wolowitz):

alright im still with u

OpenStudy (howard-wolowitz):

srry i had tot think.. but i am still with u

OpenStudy (anonymous):

ok that should be enough information to pick the correct answer

OpenStudy (howard-wolowitz):

you mean 2y

OpenStudy (anonymous):

oops, typo

OpenStudy (howard-wolowitz):

ok the first one is C .. awesome

OpenStudy (howard-wolowitz):

we got that one on LOCK DOWN

OpenStudy (anonymous):

The cost per chair A is 0.79 The total cost to produce chair A is 0.79 * x The cost per chair B is 2.00 The total cost to produce chair B is 2.00 * y total cost to produce both chairs = .79x + 2y Total cost cannot exceed 50. 0.79x + 2y <= 50

OpenStudy (howard-wolowitz):

i gotcha

OpenStudy (howard-wolowitz):

no worries

OpenStudy (anonymous):

for the next one, i dont see an easy to do it, except by trial and error

OpenStudy (howard-wolowitz):

alright i can try that

OpenStudy (howard-wolowitz):

hmm i didnt do it right